OpenStudy (anonymous):

If the length of side a is 12 centimeters, m

OpenStudy (campbell_st):

so you are dealing with a non right triangle looks like the law of sines to me you can find angle A by angle sum of a triangle A = 180 - 75 - 36 then its \[\frac{a}{\sin(A)} = \frac{b}{\sin(B)}\] hope it helps